During an earthquake the top of a high rise building sways back and forth. The largest distance between extremes of this motion are found to be 3m. A resident of the top floor reports that he timed how long it took to go from one side to the other as 1.5s.

a. Find the maximum velocity of the top floor of the building.

b. Find the maximum acceleration of the top floor of the building.
Answer #1

Given,

d = 3 m ; t = 1.5 s

a)v = d/t

v = 3/1.5 = 2 m/s

Hence, v = 2 m/s

b)a = v/t

a = 2/1.5 = 1.33 m/s^2

Hence, a = 1.33 m/s^2